OpenStudy (anonymous):

A store mixes Kenyan Coffee worth $11 per kilogram and Turkish coffee worth $15 per kilogram. The mixture is to sell for $14 per kilogram. Find how much of each should be used to make a 372-kilogram mixture. How many Kilograms of the Kenyan coffee should be in the mixture?_kg How many Kilograms of the Turkish coffee should be in the mixture?_kg

OpenStudy (mertsj):

Let x = kg of Kenyan coffee 372-x= kg of Turkish coffee 11x+15(372-x)=14(372)